OpenStudy (mitu12):

factor the polynomial w^2-9/4

OpenStudy (mitu12):

@Aureyliant

OpenStudy (mitu12):

i know it will be w+ ? and w-?

OpenStudy (anonymous):

What do you think you should do first? I literally helped you step-by-step in two prior answers so you should have some sort of understanding.

OpenStudy (mitu12):

i dont understand how to take the square root of -9/4

OpenStudy (mitu12):

i know it will be 3/2

OpenStudy (ybarrap):

Use the following rule: $$ x^2-b^2=(x+\sqrt b)(x-\sqrt b) $$

OpenStudy (mitu12):

so will the answer be (w-3/2)(w+3/2)

OpenStudy (mitu12):

@Aureyliant

OpenStudy (mitu12):

@sweetburger

OpenStudy (ybarrap):

yes!